# MARS: Unleashing the Power of Variance Reduction for Training Large Models

### v2

@@ Appendix C Proof of Auxiliary Lemmas
Appendix D Additional Experiment Results
D.1 Supplementary Results for the Main Experiments
Here we display the supplementary results for the experiments in Section 5. The training and validation losses as well as wall-clock time curves for small and medium models are displayed in Figures 6 and 7. And the 0-shot and 5-shot evaluation results on different downstream tasks for small, medium and large models are listed in Tables 2,3,4 and Tables 5,6, respectively. It can be observed that different sizes of models trained with MARS-AdamW and MARS-Lion can achieve better performances than baseline optimization methods with respect to cross-entropy loss, time efficiency as well as downstream task performances.
@@ D.1 Supplementary Results for the Main Experiments
D.2 MARS and MARS-approx
We then conduct experiments to compare the performance of MARS and MARS-approx (MARS-AdamW instantiation) on GPT-2 small and medium models, the training and validation loss curves are shown in Figures 8 and 9. Models trained with MARS exhibit consistently better performance than those trained with MARS-approx. This suggests that: (a) The exact version, which employs the variance reduction formulation, is more fundamental than the approximate version. (b) The approximate version serves as a practical alternative in scenarios where computational efficiency is a priority, as it incurs only minimal performance loss. However, in settings where maximizing validation accuracy is crucial, the exact version is recommended.
@@ D.2 MARS and MARS-approx
D.3 Experiments on FineWeb-Edu 100B Dataset
FineWeb-Edu dataset (Lozhkov et al., 2024) is a high-quality dataset based on well-filtered educational web pages. To better investigate the efficiency of our algorithm, we also use FineWeb-Edu 100B, a subset of FineWeb-Edu with around 100B tokens to train GPT-2 small (125M) and XL (1.5B, with the same learning rates as GPT-2 large models) models with optimizers including AdamW, Muon and MARS-AdamW-approx. We leave around 0.1B tokens for validation and other tokens for training. The training and evaluation curves are shown in Figures 10 and 11. It can be seen that our algorithms can also achieve better performances even with different datasets. For a comprehensive investigation, we evaluate these models on metrics same as experiences in Section 5.2, and the results are shown in Tables 7 and 8. We also compare the results with the open-source GPT-2 models on Hugging Face Radford et al. (2019) (denoted as “OpenAI-Comm.” in the tables). Compared with Table 2, it can be observed that this dataset is actually better for the superior performances. However, models trained with our algorithm can also show advantages over baseline optimization approaches trained with such a high-quality dataset.
@@ D.3 Experiments on FineWeb-Edu 100B Dataset
D.4 Computer Vision Experiments
We also carry out experiments on the classification task in the field of computer vision. We conduct experiments with ResNet-18 model (He et al., 2016) on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009) for AdamW, Lion, Shampoo [^3] In practice, we use Distributed Shampoo (Shi et al., 2023) to facilitate the training of Shampoo., Muon, and variants of MARS instantiation, following the setting in Chen et al. (2018a).
We do grid research to explore the best hyper-parameters for each of these optimization methods. We search over $\{10^{-5},...,10^{0}\}$ for the learning rate and $\{0,...,1.0\}$ for the weight decay. We set $\beta_{1}=0.9$ and search over $\{0.99,0.999\}$ for $\beta_{2}$ for AdamW and Muon; fix $\beta_{1}=0.9$ and search $\beta_{2}$ over $\{0.99,0.999\}$ for Lion; search over $\{0.9,0.95\}$ for $\beta_{1}$ and $\{0.95,0.99,0.999\}$ for $\beta_{2}$ for Shampoo; and we fix the $(\beta_{1},\beta_{2})=(0.95,0.99)$ and $\gamma=0.025$ for MARS models. We train for 200 epochs with training batch size 128 on 1 NVIDIA A6000 GPU. And we also apply MultiStepLR scheduler so that the learning rate would decrease to 10% of the original rate at the 100th epoch and to 1% at the 150th epoch. We display the test loss and test accuracy for CIFAR-10 and CIFAR-100 datasets in Figures 12 and 13, respectively. The results show that our algorithm can achieve better validation loss after the decay of learning rate and better test accuracy within the final stage of training.
@@ D.4 Computer Vision Experiments
Appendix E Hyper-parameter Settings
For training parameters, we did a grid search over learning rates between $\{1e-4,1.5e-4,3e-4,6e-4,1e-3,1.5e-3,3e-3,6e-3\}$, for weight decay coefficient, we did a grid search over $\{1e-1,1e-2,1e-3\}$. For AdamW baseline, although we utilized the golden standard learning rates in literature (a parameter search for AdamW have been done in Liu et al. (2023)), we also did a grid search on different parameters. Part of the results of different learning rates and learning rate schedules are also shown in Section 5.3.1 and 5.3.2, respectively. Table 10 summarizes the architectural hyperparameters for GPT-2 models with 125M (small), 355M (medium), 770M (large) and 1.5B (XL, only used in Appendix D.3) parameters. Table 11 lists the general hyperparameters used across all experiments, while Table 12 present the training hyperparameters for the small, medium, and large models, respectively.

### v3

@@ Appendix D Proof of Auxiliary Lemmas
Appendix E Additional Experiment Results
E.1 Supplementary Results for the Main Experiments
Here we display the supplementary results for the experiments in Section 4. The training and validation losses as well as wall-clock time curves for small and medium models are displayed in Figures 2 and 3. And the 0-shot and 5-shot evaluation results on different downstream tasks for small, medium and large models are listed in Tables 2,3,4 and Tables 5,6, respectively. It can be observed that different sizes of models trained with MARS-AdamW and MARS-Lion can achieve better performances than baseline optimization methods with respect to cross-entropy loss, time efficiency as well as downstream task performances.
@@ E.1 Supplementary Results for the Main Experiments
E.2 MARS and MARS-approx.
We then conduct experiments to compare the performance of MARS and MARS-approx (MARS-AdamW instantiation) on GPT-2 small and medium models, the training and validation loss curves are shown in Figures 4 and 5. Models trained with MARS exhibit consistently better performance than those trained with MARS-approx. This suggests that: (a) The exact version, which employs the variance reduction formulation, is more fundamental than the approximate version. (b) The approximate version serves as a practical alternative in scenarios where computational efficiency is a priority, as it incurs only minimal performance loss. However, in settings where maximizing validation accuracy is crucial, the exact version is recommended.
@@
E.3 Experiments on FineWeb-Edu 100B Dataset
FineWeb-Edu dataset (Lozhkov et al., 2024) is a high-quality dataset based on well-filtered educational web pages. To better investigate the efficiency of our algorithm, we also use FineWeb-Edu 100B, a subset of FineWeb-Edu with around 100B tokens to train GPT-2 small (125M) and XL (1.5B, with the same learning rates as GPT-2 large models) models with optimizers including AdamW, Muon and MARS-AdamW-approx. We leave around 0.1B tokens for validation and other tokens for training. The training and evaluation curves are shown in Figures 6 and 7. It can be seen that our algorithms can also achieve better performances even with different datasets. For a comprehensive investigation, we evaluate these models on metrics same as experiments in Section 4.2, and the results are shown in Tables 7 and 8. We also compare the results with the open-source GPT-2 models on Hugging Face (Radford et al., 2019) (denoted as “OpenAI-Comm.” in the tables). Compared with Table 2, it can be observed that this dataset is actually better for the superior performances. However, models trained with our algorithm can also show advantages over baseline optimization approaches trained with such a high-quality dataset.
@@ E.3 Experiments on FineWeb-Edu 100B Dataset
E.4 Computer Vision Experiments
We also carry out experiments on the classification task in the field of computer vision. We conduct experiments with ResNet-18 model (He et al., 2016) on the CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009) for AdamW, Lion, Shampoo [^1] In practice, we use Distributed Shampoo (Shi et al., 2023) to facilitate the training of Shampoo., Muon, and variants of MARS instantiation, following the setting in Chen et al. (2018).
We do grid search to explore the best hyper-parameters for each of these optimization methods. We search over $\{10^{-5},...,10^{0}\}$ for the learning rate and $\{0,...,1.0\}$ for the weight decay. We set $\beta_{1}=0.9$ and search over $\{0.99,0.999\}$ for $\beta_{2}$ for AdamW and Muon; fix $\beta_{1}=0.9$ and search $\beta_{2}$ over $\{0.99,0.999\}$ for Lion; search over $\{0.9,0.95\}$ for $\beta_{1}$ and $\{0.95,0.99,0.999\}$ for $\beta_{2}$ for Shampoo; and we fix the $(\beta_{1},\beta_{2})=(0.95,0.99)$ and $\gamma=0.025$ for MARS models. We train for 200 epochs with training batch size 128 on 1 NVIDIA A6000 GPU. And we also apply MultiStepLR scheduler so that the learning rate would decrease to 10% of the original rate at the 100th epoch and to 1% at the 150th epoch. We display the test loss and test accuracy for CIFAR-10 and CIFAR-100 datasets in Figures 8 and 9, respectively. The results show that our algorithm can achieve better validation loss after the decay of learning rate and better test accuracy within the final stage of training.
@@ E.4 Computer Vision Experiments
E.6 Different Learning Rate Scheduler
E.6.1 Constant LR
To ensure a fair comparison by eliminating the influence of learning rate changes during training and to explore the potential for continuous training with MARS, we conduct supplementary experiments on GPT-2 small, medium, and large models using a constant learning rate for both AdamW and MARS-AdamW-approx. For each group of experiments, we compared between 2 different maximum learning rates. The training and validation curves are displayed in Figures 15 and 16. The results also indicate that our algorithm has superior performances over AdamW optimizer under a fair comparison of constant learning rates.
E.6.2 WSD Scheduler
Cosine learning rate scheduler is utilized in most of our experiments. Recently Hu et al. (2024) introduced a novel learning rate scheduler called Warmup-Stable-Decay (WSD) scheduler, which composed of 3 stages, including learning rate linear-warmup stage, constant learning rate stage, as well as learning rate decay stage. To test the flexibility and ability of continuous training of MARS trained with respect to learning rate schedulers, we implement experiments on GPT-2 small and medium models with AdamW and MARS-AdamW-approx scheduled with WSD for 10k, 20k, 50k and 100k steps. The training and validation loss curves are shown in Figures 17 and 18. The curves indicate that MARS has a better potential for continuous training and exhibits an explicit edge over baseline algorithm.
E.7 Sensitivity to Batch Size
We also investigate the sensitivity to batch size of our algorithm. We implement experiments with MARS-AdamW on GPT-2 small with batch size 240, 480 or 960, and compare them to AdamW. We fix other hyper-parameters the same as the main experiments and trained with 80,000 steps. The results are plotted in Figure 19. It can be seen that MARS-AdamW consistently outperforms AdamW, with the performance gap widening at smaller batch sizes—supporting the intuition that variance reduction is especially effective in high-variance (small-batch) settings.
Figure 19: The training and validation loss curves for varying global batch sizes (240/480/960), plotted against iteration steps on GPT-2 small model (125M).
Appendix F Hyper-parameter Settings
For training parameters, we did a grid search over learning rates between $\{1e-4,1.5e-4,3e-4,6e-4,1e-3,1.5e-3,3e-3,6e-3\}$, for weight decay coefficient, we did a grid search over $\{1e-1,1e-2,1e-3\}$. For AdamW baseline, although we utilized the golden standard learning rates in literature (a parameter search for AdamW have been done in Liu et al. (2023)), we also did a grid search on different parameters. Part of the results of different learning rates and learning rate schedules are also shown in Section E.6.1 and E.6.2, respectively. Table 10 summarizes the architectural hyperparameters for GPT-2 models with 125M (small), 355M (medium), 770M (large) and 1.5B (XL, only used in Appendix E.3) parameters. Table 11 lists the general hyperparameters used across all experiments, while Table 12 present the training hyperparameters for the small, medium, and large models, respectively.